# Approximation and decomposition of attractors of a Hopfield neural network system

### v1

@@ 4 PS algorithm for attractors of the HNN system
To facilitate the understanding of attractors approximation and decomposition of the HNN system, one can evaluate the bifurcation diagrams within a small neighborhood of $p=1$ (Fig. 3), generated with the same RK4 scheme as used for the PS algorithm. The two colors, red and blue, indicate two symmetric with respect to the origin determined by the two equilibria.
@@
As shown in [7] (where $p=1$), for the coefficients given by $W$ in (3), the HNN system has only chaotic transients, not attractors, which are proven to be hidden too [7]. Thus, after a relatively long integration time the chaotic behavior of these two transients (denoted $T_{1,2}$ in Fig. 4 (a), light red and plot respectively) vanishes and the trajectories reach one of the hidden stable cycles $C_{1,2}$ (red and blue thick plots in Fig. 4 (a)). From the time series and zooming in Fig. 4 (b), one deduces that the chaotic transient lasts till, about $t=1300$. To note that if for $p=1$, one use the matlab routine $ode45$ (utilized in several works on HNN systems), the obtained trajectories generate apparently chaotic attractors, which seems to verify numerically the asymptotic definition of attractors ($T_{1}^{\prime}$ and $T_{2}^{\prime}$ in Figs. 4 (c), (d)). However, increasing the precision of the routine $ode45$ by using e.g. ’$options=odeset(^{\prime}RelTol^{\prime},1e-6,^{\prime}AbsTol^{\prime},1e-9)$’, the trajectories reveal the behavior of the chaotic transients presented in Figs. 4 (a), (b), as obtained by the RK4 method.
The bifurcation diagram together with intensive numerical simulations suggest the following result (see [7], where the HNN is implemented for $p=1$)
@@
For $p=1$, the HNN system with the coefficients (3) admit only chaotic transients sets, no chaotic attractors.
@@ 4.1 Approximation of HNN attractors
Chaotic dynamics in neuronal systems might be useful, and even necessary. On the other side, there are situations when the emerging chaotic dynamics are undesirable [29]. In both situations, the PS algorithm could be use to give a satisfactory solution.
Consider the plausible situation when, by some objective reasons, the underlying parameter value of some desired attractor of the HNN system (4) cannot be set thereby the numerical integration cannot lead to that attractor. In these situations, via Theorem 3, Corollary 6 and Corollary 8, the PS algorithm provides the following result
@@
For example, to “force” the system to evolve e.g. along the stable cycle $A_{1.004}$ (see Fig. 5) corresponding to $p=1.004$ (assuming this value cannot be set directly), one can consider the following scheme:
@@
The non-uniqueness of the scheme $S$ to approximate a desired attractor (e,g, schemes (10) and (11)), is important because offers the possibility to choose among several variants to generate an approximation of the considered attractor.
@@
On the other side, if one switches the parameter $p$ within a set $\mathcal{P}_{N}$, which generates regular dynamics, it is possible to find the weights $\mathcal{M}_{N}$ such that the obtained attractor is chaotic (anticontrol-like action). For example, if one uses the set of parameters which generate the stable cycles $A_{1}$, $A_{1.004}$ and $A_{1.0072}$ (see the bifurcation diagram in Fig. 4), with the scheme
@@
one obtains the chaotic attractor $A^{*}$, approximation of the attractor $A_{1.0051}$.
@@ 4.2 HNN attractors as convex combinations of other attractors
which suggests that a chaotic attractor can be decomposed function of regular attractors.
Moreover, a regular attractor, as $A_{1.004}$, can be expressed as combination of other regular attractors (such as the attractor $A_{1.004}$ obtained from the regular attractors $A_{1}$, $A_{1.072}$). Similarly, chaotic attractors can be decomposed as combination of other chaotic attractors.
@@
In Appendix a short matlab code $PS.m$ is presented (a modified variant of the code presented in [6]), runs the PS algorithm for a larger number $N$ of parameters, chosen randomly.
The code allows to verify that even for random values of $p$ and $m$ (chosen with respect the algorithm requirements), the obtained results still sustain the analytical results. The code uses the function $rand$, which generates real numbers with uniform distribution in the interval $(0,1)$.
@@
The program asks the integration time $T_{max}$, the integration step size $h$, $N$, the interval $[a,b]$ within $p$ is randomly generated, the largest value of $m$, $Mmax$, $m<Mmax$, and initial conditions $x_{0}$ and $y_{0}$. The outputs are the set $\mathcal{P}$, $\mathcal{M}$, $p_{0}$, and the averaged solution $x$ and the switched solution $y$.
@@
which generates 50 random parameter values within $[0.999,1.008]$ (corresponding to the bifurcation diagram in Fig. 3), with random weights $m<7$, the integration step size $h=0.005$, and initial conditions $[0.1,0.1,0.1]$ and $[0.2,0.2,0.2]$, one obtains the chaotic attractors in Fig. 6 (a), where the switched and average attractors are overplotted. The generated values of $p$ and $m$ are presented in Table (a) (with first $8$ decimals displayed), which generates
@@ 5 Conclusions and Discussion
In this paper, the PS algorithm is briefly introduced and is applied for the first time to approximate and decompose attractors of the HNN system (4). For the great majority of continuous chaotic systems, such as Lorenz, Roessler, Chen systems etc., a chaotic system can be brought to the form required by the PS algorithm (1)-(2), i.e. a linear dependence of a bifurcation parameter, $p$. It is underlined that, for the bifurcation parameter $p=1$, with the coefficients $W$ given by (3), the HNN system has only chaotic transient sets, not chaotic attractors. Also, by using the opportunity given by the PS algorithm, to decompose an attractor as a convex combination of a set of other given attractors, it is shown both analytically and numerically too that every attractor of the HNN system can be considered as a convex combination of other attractors. Moreover, it is shown that this decomposition can be interpreted as a control-like or anticontrol-like technique: the PS algorithm allows to approximate a regular attractor or a chaotic attractor, respectively. The used numerical scheme to integrate the IVP is, RK4. It is warned that, without $options$, the matlab routine $ode45$ could give for this system wrong results concerning the chaotic attractors.
The authors declare that they have no conflict of interest.

### v2

@@ 4 PS algorithm for attractors of the HNN system
To facilitate the understanding of attractors approximation and decomposition of the HNN system, one can evaluate the bifurcation diagrams within a small neighborhood of $p=1$ (Fig. 3), generated with the same RK4 scheme as used for the PS algorithm. The two colors, red and blue, indicate two symmetric with respect to the origin determined by the two equilibria. The initial conditions for the switched solution, and for the average solution, $x_{0}$ and $y_{0}$, respectively, are $(0.1,0.1,0.1)$ and $(0.2,0.2,.2)$.
@@
As shown in [5] (where $p=1$), for the coefficients given by $W$ in (3), the HNN system has only chaotic transients, not attractors, which are proven to be hidden too [5]. Thus, after a relatively long integration time the chaotic behavior of these two transients (denoted $T_{1,2}$ in Fig. 4 (a), light red, and plot respectively) vanishes, and the trajectories reach one of the hidden stable cycles $C_{1,2}$ (red and blue thick plots in Fig. 4 (a)). From the time series and zooming in Fig. 4 (b), one deduces that the chaotic transient lasts till, about $t=1300$. Note that if for $p=1$, one use the Matlab routine $ode45$ (utilized in several works on HNN systems), the obtained trajectories generate apparently chaotic attractors, which seems to verify numerically the asymptotic definition of attractors ($T_{1}^{\prime}$ and $T_{2}^{\prime}$ in Figs. 4 (c), (d)). However, increasing the precision of the routine $ode45$ by using e.g. ’$options=odeset(^{\prime}RelTol^{\prime},1e-6,^{\prime}AbsTol^{\prime},1e-9)$’, the trajectories reveal the behavior of the chaotic transients presented in Figs. 4 (a), (b), as obtained by the RK4 method.
The bifurcation diagram together, with intensive numerical simulations, suggest the following result (see [5], where the HNN is implemented for $p=1$)
@@
For $p=1$, the HNN system with the coefficients (3) admits only chaotic transients sets, no chaotic attractors.
@@ 4.1 Approximation of HNN attractors
Chaotic dynamics in neuronal systems might be useful, and even necessary. On the other side, there are situations when the emerging chaotic dynamics are undesirable [36]. In both situations, the PS algorithm could be used to give a satisfactory solution.
Consider the plausible situation when, for some objective reasons, the underlying parameter value of some desired attractor of the HNN system (4) cannot be set thereby the numerical integration cannot lead to that attractor. In these situations, via Theorem 3, Corollary 6 and Corollary 8, the PS algorithm provides the following result
@@
For example, to “force” the system to evolve, e.g. along the stable cycle $A_{1.004}$ (see Fig. 5) corresponding to $p=1.004$ (assuming this value cannot be set directly), one can consider the following scheme:
@@
The non-uniqueness of the scheme $S$ to approximate a desired attractor (e,g, schemes (10) and (11)), is important because it offers the possibility to choose among several variants to generate an approximation of the considered attractor.
@@
On the other side, if one switches the parameter $p$ within a set $\mathcal{P}_{N}$, which generates regular dynamics, it is possible to find the weights $\mathcal{M}_{N}$ such that the obtained attractor is chaotic (anticontrol-like action). For example, if one uses the set of parameters that generate the stable cycles $A_{1}$, $A_{1.004}$ and $A_{1.0072}$ (see the bifurcation diagram in Fig. 4), with the scheme
@@
one obtains the chaotic attractor $A^{*}$, an approximation of the attractor $A_{1.0051}$.
@@ 4.2 HNN attractors as convex combinations of other attractors
which suggests that a chaotic attractor can be decomposed using regular attractors.
Moreover, a regular attractor, as $A_{1.004}$, can be expressed as combination of other regular attractors (such as the attractor $A_{1.004}$ obtained from the regular attractors $A_{1}$, $A_{1.072}$). Similarly, chaotic attractors can be decomposed as a combination of other chaotic attractors.
@@
In Appendix, a short Matlab code $PS.m$ is presented (a modified variant of the code presented in [1]), runs the PS algorithm for a larger number $N$ of parameters, chosen randomly.
The code allows to verify that even for random values of $p$ and $m$ (chosen with respect to the algorithm requirements), the obtained results still sustain the analytical results. The code uses the function $rand$, which generates real numbers with a uniform distribution in the interval $(0,1)$.
@@
The program asks the integration time $T_{max}$, the integration step size $h$, $N$, the interval $[a,b]$ within $p$ is randomly generated, the largest value of $m$, $Mmax$, $m<Mmax$, and the initial conditions $x_{0}$ and $y_{0}$. The outputs are the set $\mathcal{P}$, $\mathcal{M}$, $p_{0}$, and the averaged solution $x$ and the switched solution $y$.
@@
which generates 50 random parameter values within $[0.999,1.008]$ (corresponding to the bifurcation diagram in Fig. 3), with random weights $m<7$, the integration step size $h=0.005$, and initial conditions $[0.1,0.1,0.1]$ and $[0.2,0.2,0.2]$, one obtains the chaotic attractors in Fig. 6 (a), where the switched and average attractors are overplotted. The generated values of $p$ and $m$ are presented in Table (a) (with the first $8$ decimals displayed), which generates
@@ 5 Conclusions and Discussion
In this paper, the PS algorithm is briefly introduced and is applied for the first time to approximate and decompose attractors of the HNN system (4). For the great majority of continuous chaotic systems, such as Lorenz, Roessler, Chen system etc., a chaotic system can be brought to the form required by the PS algorithm (1)-(2), i.e., a linear dependence of a bifurcation parameter, $p$. It is underlined that, for the bifurcation parameter $p=1$, with the coefficients $W$ given by (3), the HNN system has only chaotic transient sets, not chaotic attractors. Also, by using the opportunity given by the PS algorithm, to decompose an attractor as a convex combination of a set of other given attractors, it is shown both analytically and numerically too, that every attractor of the HNN system can be considered as a convex combination of other attractors. Moreover, it is shown that this decomposition can be interpreted as a control-like or anticontrol-like technique: the PS algorithm allows to approximate a regular attractor or a chaotic attractor, respectively. The numerical scheme used to integrate the IVP is, RK4. It is warned that, without $options$, the Matlab routine $ode45$ could give for this system wrong results concerning the chaotic attractors. A future study will be carried out to analyze the potential algebraic structure defined by the two binary operations $\oplus$ and $\odot$ over the set of attractors of a dynamical system modeled by (1)-(2).